OpenStudy (anonymous):

classify the conic section and write its equation in standard form: #1- 9x^2+49y^2+392y+343=0 #2- x^2+y^2+4y-5=0

myininaya (myininaya):

If you have the signs in front of x^2 and y^2 have different signs, then most likely you have a hyperbola it's standard form is \[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1 \text{ or } \frac{(y-k)^2}{b^2}-\frac{(x-h)^2}{a^2}=1 \] If you have the same sign but the coefficients are not equal (still talking about x^2 and y^2), then most likely you have an ellipse it's standard form is \[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2 }{b^2}=1 \] if the coefficients are the same (still talking about x^2 and y^2), then most likely it is a circle it's standard form is \[(x-h)^2+(y-k)^2=r^2 \] Now a parabola, one of the coefficients of x^2 or y^2 will be 0 (but not both) and it's standard form is \[(x-h)^2=4p(y-k) \text{ or } (y-k)^2=4p(x-h) \]

myininaya (myininaya):

so out of what I just say, your equations are most likely what?

OpenStudy (anonymous):

the answer choices for #1 are A.) hyperbola (y-1)^2/4-x^2/25=1 B.) ellipse x^2/49+(y+4)^2/9=1 C.) ellipse x^2/9+(y+4)^2/49=1 D.)parabola x=-2y^2=3

myininaya (myininaya):

ok and I'm asking you to look at what I said and tell what kinda graph it is most likely to give us

OpenStudy (anonymous):

Oh, sorry I thought you were asking me what my equations were, read that wrong sorry. Okay so would it be a...hyperbola? because the signs in front of y^2 and x^2 are different..

myininaya (myininaya):

so that - thing in front of your equation is part of your problem? like it is a negative sign or just the way they are numbered 1- 2-

OpenStudy (anonymous):

that's just the way they are numbered 1- 2- and then the equation

myininaya (myininaya):

so that would mean the x^2 and y^2 coefficients are both same in sign then

myininaya (myininaya):

so definitely not a hyperbola

myininaya (myininaya):

hyperbola you would have different signs

OpenStudy (anonymous):

oh I thought you meant how ones 9x^2 and ones 49y^2, so then would it be a ellipse right?